OpenStudy (anonymous):

Find the equation of the ellipse Center at (2,0), Focus (2,6), Eccentricity 2/3. Can someone help me please? Thanks. :)

OpenStudy (anonymous):

oops, I blew it?

OpenStudy (amistre64):

eccentricity is one that never came up i class ... i know it has something to do with a and b tho

OpenStudy (amistre64):

c/a is what the interweb tells me

OpenStudy (anonymous):

e = sqrt(1- a^2/b^2) Am wondering if I have to flip them because ellipse is rotated. Might have to go back to basics, bring back to to (0,0) calculate then retranslate rotate...yuk!

OpenStudy (amistre64):

c is your focal length and a is your vertex length; and b^2 = a^2-c^2 to find the short vertex

OpenStudy (amistre64):

rotate? no; just x^2 y^2 --- + --- = 1 a^2 b^2 a and b can change depending on the major axis

OpenStudy (anonymous):

e = sqrt(1- a^2/b^2) Am wondering if I have to flip them because ellipse is rotated. Might have to go back to basics, bring back to to (0,0) calculate then retranslate rotate...yuk!

OpenStudy (anonymous):

The llipse she is describing is rotated on end, so you have to swap a, b.

OpenStudy (amistre64):

since this is elongated along the y axis; a goes under y

OpenStudy (anonymous):

I answered it once, guess I got it wrong though.

OpenStudy (anonymous):

e = sqrt(1- a^2/b^2) Am wondering if I have to flip them because ellipse is rotated. Might have to go back to basics, bring back to to (0,0) calculate then retranslate rotate...yuk!

OpenStudy (anonymous):

The basic equation isn't a problem, it's actually calculating the a and b with eccentricity and focus info.

OpenStudy (amistre64):

Center (2,0), Focus (2,6), focal length = 6, which simplifies Eccentricity 2(3)/3(3) = 6/9 (x-2)^2 (y)^2 ------ + ----- = 1 b^2 81 b^2 = 9^2-6^2 = 81-36 = 45 right?